OpenStudy (anonymous):

Find 16,5^(1/4) without using computer.

OpenStudy (anonymous):

Calculator I meant,

OpenStudy (solomonzelman):

\[\huge\color{green}{ a^ {(b/c)}=\sqrt[c]{a^b} } \]

OpenStudy (solomonzelman):

\[\huge\color{green}{ 16.5^ {(1/4)}=\sqrt[4]{16.5^1}=\sqrt[4]{16.5} } \]

OpenStudy (solomonzelman):

This can be simplified more though, can you simplify this root?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):